OpenStudy (anonymous):

How many three symbol codes using the digits 1 through 9 are possible if no repetitions are allowed?

jimthompson5910 (jim_thompson5910):

There are 9 digits, so there are 9*8*7 = 72*7 = 504 different possible codes (where order matters and repetition is not allowed)